# How Hyper-Datafication Impacts the Sustainability Costs in Frontier AI

### v2

@@ Abstract.
Large-scale data has fuelled the success of frontier artificial intelligence (AI) models over the past decade. This expansion has relied on sustained efforts by large technology corporations to aggregate and curate internet-scale datasets. In this work, we examine the environmental, social, and economic costs of large-scale data in AI through a sustainability lens. We argue that the field is shifting from building models from data to actively creating data for building models. We characterise this transition as hyper-datafication, which marks a critical juncture for the future of frontier AI and its societal impacts. To quantify and contextualise data-related costs, we analyse approximately 550,000 datasets from the Hugging Face Hub, focusing on dataset growth, storage-related energy consumption and carbon footprint, and societal representation using language data. We complement this analysis with qualitative responses from data workers in Kenya to examine the labour involved, including direct employment by big tech corporations and exposure to graphic content. We further draw on external data sources to substantiate our findings by illustrating the global disparity in data centre infrastructure. Our analyses reveal that hyper-datafication does not merely increase resource consumption but systematically redistributes environmental burdens, labour risks, and representational harms toward the Global South, precarious data workers, and under-represented cultures. Thus, we propose Data PROOFS recommendations spanning provenance, resource awareness, ownership, openness, frugality, and standards to mitigate these costs. Our work aims to make visible the often-overlooked costs of data that underpin frontier AI and to stimulate broader debate within the research community and beyond. [^1] Source code used to extract metadata from datasets hosted on the Hugging Face Hub is available at: https://github.com/saintslab/costs-of-hyperdatafication.
sustainability, data, artificial intelligence, social impact, carbon footprint
copyright: nonejournalyear: 2026ccs: Computing methodologies Artificial intelligenceccs: Hardware Impact on the environment
@@ 1. Introduction
These investments would be futile without access to another critical resource: data. Contemporary frontier AI models train on datasets containing tens of trillions of data points, often scraped at scale from the internet. For example, one of the largest publicly available tokenised datasets DCLM-Pool, contains more than 240 trillion tokens (Li et al., 2024). Figure 1 illustrates the rapid growth of datasets and data volume using the datasets available on the Hugging Face Hub [^2] https://huggingface.co/datasets, a platform for hosting AI models and datasets (see Section 3).
@@
Growing concern about the sustainability of AI has prompted an expanding body of work (Van Wynsberghe, 2021; Wu et al., 2022; Selvan, 2025). However, existing discussions largely centre on model training and deployment. The sustainability of data itself remains comparatively under-examined, despite the fact that collecting and using internet-scale datasets for frontier AI is highly resource-intensive and can generate substantial environmental, social, and economic costs. As a result, current sustainability discussions overlook data as a central driver of AI’s broader impacts.
@@
In this work, we take a comprehensive look at the sustainability of data for AI. Building on existing critiques of the sustainability of AI (Van Wynsberghe, 2021; Wu et al., 2022; Wright et al., 2023), we focus explicitly on the environmental, social, and economic costs of large-scale data practices. This perspective becomes increasingly urgent as frontier AI research pursues ever more ambitious goals, including artificial general intelligence (AGI), intensifying the demand for data at scale.
@@
We argue that the field is undergoing a structural shift in how data is produced and used. First, frontier AI systems increasingly rely on large-scale synthetic data. For example, the Phi-4 model was pretrained on approximately 10 trillion tokens, of which around 40% were synthetically generated and curated using other AI models (Abdin et al., 2024). Second, and more fundamentally, AI development is driving the active creation of new data sources designed explicitly for model training. A prominent example is the Ego4D dataset, which consists of 3,670 hours of egocentric video collected from 923 participants across 74 locations, recorded specifically to support AI systems (Grauman et al., 2022). Such datasets would not exist outside the demands of model development.
@@
We use the term hyper-datafication to characterise this convergence, which we define formally below. While the term has appeared sporadically in prior work, it has been used narrowly to describe the large-scale expansion and aggregation of existing data sources (Rosales and Suárez-Gonzalo, 2024). We extend this usage by explicitly incorporating two additional, increasingly central dimensions: the algorithmic generation of synthetic data and the creation of data whose primary purpose is to serve as training material for AI models rather than to support direct human activities.
@@
Hyper-datafication refers to the industrialised production and accumulation of data for AI model development across three coupled processes: (i) the large-scale scaling and recombination of existing data sources, (ii) the use of AI systems to generate synthetic data, and (iii) the creation of purpose-built data whose primary function is to serve as training input for AI systems rather than direct human use.
We analyse the costs of hyper-datafication across the different sustainability dimensions. To support this analysis, we draw on multiple empirical sources. First, we conduct a large-scale metadata study of approximately $550,000$ datasets hosted on the Hugging Face Hub, documenting growth in dataset scale, size, and scope. Using this metadata, we estimate storage-related energy consumption and associated carbon footprints on both the provider and user side, providing indicators of broader environmental trends. We then first examine social costs through a labour-focused analysis. Here, we draw on qualitative evidence from a structured questionnaire administered to data workers in Kenya to illustrate how hyper-datafication shapes the scale, nature, and conditions of contemporary data work. We further examine social costs by analysing patterns of language representation in the Hugging Face Datasets, assessing how supply and demand for datasets are distributed across languages relative to global speaker populations and web presence. This analysis highlights systematic linguistic concentration and under-representation, revealing how hyper-datafication amplifies existing asymmetries in whose languages and cultures are most extensively captured in AI training data. Finally, we incorporate external data sources to contextualise and substantiate our economic analysis and to support claims in the environmental and social sections, drawing on reports, policy documents, and prior empirical studies where direct measurement is not feasible. Based on these analyses, we propose actionable recommendations, formulated as the Data PROOFS guidelines, spanning provenance, resource awareness, ownership, openness, frugality, and standards to mitigate the costs of hyper-datafication.
@@
The following sections briefly discuss existing research on the sustainability impacts of AI systems, and the importance of data in this context (Section 2). We describe our methods for gathering information on the sustainability implications of AI-related data (Section 3), and we analyse our findings in light of the growing trend of hyper-datafication (Section 4). We conclude with a set of recommendations (Data PROOFS) to increase data sustainability and address the actions required to reduce the negative impacts of hyper-datafication (Section 5).
@@ 2.1. Sustainability of AI
Broader Environmental Impact of AI. Data centres use vast volumes of fresh water for cooling (Li et al., 2025) threatening freshwater resources in vulnerable regions (Environmental and Energy Study Institute, 2024). The manufacture of hardware, use of rare earths, electronic waste generation, and land-use change associated with AI data infrastructure contribute major embodied carbon emissions and ecological damage (Kaack et al., 2022; Wright et al., 2023). The amount of e-waste has grown from 34 million tons in 2010 to 64 million tons in 2022, and is projected to grow to 82 million tons by 2030. During the same time, the fraction of e-waste recycled decreased from 24% to less than 5% (UNITAR et al., 2024).
@@ 2.2. Costs of Data
Environmental. While there are no comprehensive studies that estimate the environmental cost of AI datasets, there are specialised studies that focus on domain-specific data. For example, Souter et al. (2025) study the impact of data preprocessing in medical image analysis, benchmarking the effects of different preprocessing stages. In Wang et al., authors estimate the direct and embodied emissions due to global land cover mapping projects which is a data-intensive task due to the high-resolution satellite images (Wang et al., 2025). They estimated the emissions between 2014 and 2024 to be about 3.84 million tCO2eq, and project it to increase 60 times to about 184 million tCO2eq by 2050. Outside of AI, carbon emissions across the data lifecycle have been explored by Mersy and Krishnan (2024); they propose carbon provenance to annotate data with carbon emission meta-data, facilitating better carbon accounting of data.
Economic. Research on the economic consequences of AI data highlights the uneven distribution of the burdens and the benefits across regions and actors. Studies show that AI development reproduces structural inequalities between the Minority World and the Majority World (Mohamed et al., 2020; Trabelsi, 2024). The extreme market concentration reinforces this inequality: a small group of dominant platforms holds disproportionate control over behavioural and operational data, creating data monopolies and high barriers to entry for smaller actors (Zuboff, 2023; Schweitzer et al., 2019). By analysing the economic data value chain, Jia et al. (2025) argue that monetary value systematically accrues to aggregators and model developers, i.e., the dominant platforms, while data generators are largely excluded.
@@ 3. Methodology
To address the costs of hyper-datafication holistically, we estimate the energy use of large-scale data storage (environmental), analyse labour conditions using questionnaires administered to data workers and patterns of language representation in AI datasets (social), and draw on public data to examine economic costs associated with AI data (economical). We aim to identify: (i) the environmental impacts of data storage under hyper-datafication; (ii) the actors and labour conditions that sustain hyper-datafication; (iii) the extent to which hyper-datafication represents society using language as a proxy; and (iv) the economic burdens associated with hyper-datafication.
@@ 3.1. Metadata of Hugging Face Datasets
Scope. We retrieved metadata on 1 December 2025, when the Hugging Face Hub hosted 570,802 datasets. Metadata extraction failed for 2.89% of repositories, yielding a final sample of 554,300 datasets. We collected repository-level metadata (identifiers, timestamps, downloads, and Hub-side storage), dataset-level metadata (dataset size), and contextual attributes (region, modality, task, and language). Hugging Face has options to query download statistics in two granularities: all time and last 30 days. As we retrieved our data on 1 December 2025, the statistics for last 30 days correspond to November 2025.
@@
Data collection. We collected repository-level metadata via the Hugging Face REST API [^3] Using the endpoint https://huggingface.co/api/datasets/{id} (one request per dataset ID) and retrieving attributes via the API’s expand fields., obtained dataset size statistics separately from the Hugging Face datasets-server API [^4] See https://github.com/huggingface/dataset-viewer for documentation of the datasets-server infrastructure., and extracted contextual attributes from the Hugging Face Hub Python API [^5] Accessing the public dataset catalogue at https://huggingface.co/datasets by querying the full public dataset catalogue. These attributes are primarily self-declared and therefore exhibit lower coverage. Appendix A.1 summarises all metadata attributes and coverage.
@@
Data analysis. We performed basic preprocessing and used the dataset size as a proxy for local storage footprint, while Hub-side storage was used as a proxy for platform storage. We mapped language tags to ISO-639 codes. For the language analysis, dataset volumes were compared against global speaker distributions and web presence, using Common Crawl page counts as a proxy for online content. Extended descriptive statistics are provided in Appendix A.2.
3.2. Questionnaire for Data Workers
@@
To assess social and labour costs, we conducted an online questionnaire targeting data workers in Kenya.
Participants. The final sample includes 134 respondents located in Kenya. The sample comprises 57 females and 77 males, predominantly aged 20-40 years, with most reporting at least four years of experience. Respondents were informed about the study’s purpose and how their data would be used. They were free to participate and stop the survey at any time and were unpaid.
@@
Data collection. We administered an English-language questionnaire via an online form and distributed it via the messaging application Telegram to data workers who had voluntarily joined the Telegram channel organised by a local data workers’ collective. Data collection took place in December 2025. The questionnaire comprised ten questions covering demographics, working conditions, exposure to graphic content, and employment relationships with large technology companies. The complete questionnaire is provided in Appendix B.
@@
Data analysis. We computed descriptive statistics and pairwise associations across working hours, experience, salary, exposure to graphic content, and employment type. Some questions allowed free-text responses, which, in a few cases, were used to harmonise categories. Free-text responses that could not be mapped consistently to the catergories were excluded. Detailed preprocessing decisions and complete raw counts are reported in Appendix C.

### v3

@@ Abstract.
Large-scale data has fuelled the success of frontier artificial intelligence (AI) models over the past decade. This expansion has relied on sustained efforts by large technology corporations to aggregate and curate internet-scale datasets. In this work, we examine the environmental, social, and economic costs of large-scale data in AI through a sustainability lens. We argue that the field is shifting from building models from data to actively creating data for building models. We characterise this transition as hyper-datafication, which marks a critical juncture for the future of frontier AI and its societal impacts. To quantify and contextualise data-related costs, we analyse approximately 550,000 datasets from the Hugging Face Hub, focusing on dataset growth, storage-related energy consumption and carbon footprint, and societal representation using language data. We complement this analysis with qualitative responses from data workers in Kenya to examine the labour involved, including direct employment by big tech corporations and exposure to graphic content. We further draw on external data sources to substantiate our findings by illustrating the global disparity in data centre infrastructure. Our analyses reveal that hyper-datafication does not merely increase resource consumption but systematically redistributes environmental burdens, labour risks, and representational harms toward the Global South, precarious data workers, and under-represented cultures. Thus, we propose Data PROOFS recommendations spanning provenance, resource awareness, ownership, openness, frugality, and standards to mitigate these costs. Our work aims to make visible the often-overlooked costs of data that underpin frontier AI and to stimulate broader debate within the research community and beyond.
sustainability, data, artificial intelligence, social impact, carbon footprint, data labour
copyright: noneccs: Computing methodologies Artificial intelligenceccs: Hardware Impact on the environment
@@ 1. Introduction
These investments would be futile without access to another critical resource: data. Contemporary frontier AI models train on datasets containing hundreds of trillions of data points often scraped at scale from the internet. For example, one of the largest publicly available tokenised datasets DCLM-Pool, contains more than 240 trillion tokens (Li et al., 2024). Figure 1 illustrates the rapid growth of datasets and data volume using the datasets available on the Hugging Face Hub [^1] https://huggingface.co/datasets, a public platform for hosting AI models and datasets (see Section 3).
@@
Growing concern about the sustainability of AI has prompted an expanding body of work (Van Wynsberghe, 2021; Wu et al., 2022; Selvan, 2025). However, existing discussions typically focus on a single dimension of sustainability and largely centre on model training and deployment. The sustainability of data itself remains comparatively under-examined, despite the fact that collecting and using internet-scale datasets for frontier AI is highly resource-intensive and can generate substantial environmental, social, and economic costs. As a result, current sustainability discussions overlook data as a central driver of AI’s broader impacts.
@@
In this work, we take a comprehensive look at the sustainability of data for AI. Building on existing critiques of the sustainability of AI (Van Wynsberghe, 2021; Wu et al., 2022; Wright et al., 2023), we shift the focus from models to data and explicitly consider the environmental, social, and economic costs together. This joint perspective reflects that these costs are inherently interdependent and must be understood within the broader context of holistic sustainability. Optimising one dimension in isolation risks shifting burdens elsewhere. This perspective becomes increasingly urgent as frontier AI research pursues ever more ambitious goals, including artificial general intelligence (AGI), intensifying the demand for data at scale.
At the same time, however, access to large-scale data has enabled substantial societal benefits. In climate modelling, data-driven systems can now predict climate-related disasters such as floods and wildfires with high accuracy (Jailani and Nurmadewi, 2025; Papakis et al., 2025), improving early warning systems and disaster preparedness. In health care, large-scale data has enabled breakthroughs in drug-discovery, including protein-folding (Jumper et al., 2021), supports earlier disease detection (Litjens et al., 2016; Kooi et al., 2017; Lauritzen et al., 2022) and more personalised treatment strategies (Fung et al., 2020; Gu et al., 2022). Such applications demonstrate the transformative potential of large-scale data to directly save and improve human lives.
@@
While these benefits are real and significant, we argue that the field is undergoing a structural shift in how data is produced and used. Increasingly, large-scale data is not only leveraged to address concrete societal challenges, but is created primarily to sustain model scaling. One manifestation of this shift is the growing reliance on synthetic data in frontier AI systems. For example, the Phi-4 model was pretrained on approximately 10 trillion tokens, of which around 40% were synthetically generated and curated using other AI models (Abdin et al., 2024). More fundamentally, AI development is driving the active creation of new data sources designed explicitly for model training. A prominent example is the Ego4D dataset, which consists of 3,670 hours of egocentric video collected from 923 participants across 74 locations, recorded specifically to support AI systems (Grauman et al., 2022). Such datasets would not exist outside the demands of model development.
@@
We use the term hyper-datafication to characterise these trends, which we define formally below. While the term has appeared sporadically in prior work, it has been used narrowly to describe the large-scale expansion and aggregation of existing data sources (Rosales and Suárez-Gonzalo, 2024). We extend this usage by explicitly incorporating two additional, increasingly central dimensions: the algorithmic generation of synthetic data and the creation of data whose primary purpose is to serve as training material for AI models rather than to support direct human activities.
@@
Hyper-datafication refers to the industrialised production and accumulation of data for AI model development across three coupled processes: (i) the large-scale data collection and recombination of existing data sources, (ii) the use of AI systems to generate synthetic data, and (iii) the creation of purpose-built data whose primary function is to serve as training input for AI systems rather than direct human use.
Assessing the sustainability costs of hyper-datafication requires examining how large-scale data production, storage, and downstream use translate into environmental, social, and economic costs. Despite the central role of data in frontier AI systems, these dimensions remain only partially documented (see Section 2). To analyse these costs, we draw on multiple complementary sources: a large-scale metadata study of datasets stored on the Hugging Face Hub; empirical evidence from data workers in Kenya; and external reports and public data to contextualise global data-centre investments and infrastructure. The main contributions of this work are:
  • A large-scale empirical analysis of more than 550,000 datasets stored on the Hugging Face Hub, documenting growth trends, storage- and carbon-related energy consumption for both provider and user sides, and patterns of linguistic representation relative to global speaker populations and web presence.
  • Qualitative evidence from 134 data workers in Kenya, revealing conditions of low pay, lack of compensation for exposure to graphic content, direct employment relationships with large technology companies, and gender-disaggregated disparities.
  • An analysis of the global investments and geographical distribution of data-centres, drawing on external data sources and public reports to document disparities in data centre infrastructure investment that concentrate economic power among a small number of countries.
  • The Data PROOFS recommendations, a framework outlining actions required to increase data sustainability and reduce the negative impacts of hyper-datafication.
@@
The following section briefly discusses existing research on the sustainability impacts of AI systems and the importance of data in this context (Section 2). We then describe our methods for gathering and analysing information on the sustainability implications of AI-related data (Section 3), before presenting our findings in light of the growing trend of hyper-datafication (Section 4). We conclude with a set of recommendations under the Data PROOFS framework (Section 5).
@@ 2.1. Sustainability of AI
Broader Environmental Impact of AI. Data centres use vast volumes of fresh water for cooling (Li et al., 2025) threatening freshwater resources in vulnerable regions (Environmental and Energy Study Institute, 2024). The manufacture of hardware, use of rare earths, electronic waste (e-waste) generation, and land-use change associated with AI data infrastructure contribute major embodied carbon emissions and ecological damage (Kaack et al., 2022; Wright et al., 2023). The amount of e-waste has grown from 34 million tons in 2010 to 64 million tons in 2022, and is projected to grow to 82 million tons by 2030. During the same time, the fraction of e-waste recycled decreased from 24% to less than 5% (UNITAR et al., 2024).
@@ 2.2. Costs of Data
Environmental. While there are no comprehensive studies that estimate the environmental cost of AI datasets, there are specialised studies that focus on domain-specific data. For example, Souter et al. (2025) study the impact of data preprocessing in medical image analysis, benchmarking the effects of different preprocessing stages. In Wang et al., authors estimate the direct and embodied emissions due to global land cover mapping projects which is a data-intensive task due to the high-resolution satellite images (Wang et al., 2025). They estimated the emissions between 2014 and 2024 to be about 3.89 million tCO2eq, and project it to increase to 184 million tCO2eq by 2050. Outside of AI, carbon emissions across the data lifecycle have been explored by Mersy and Krishnan (2024); they propose carbon provenance to annotate data with carbon emission meta-data, facilitating better carbon accounting of data.
Economic. Research on the economic consequences of AI data highlights the uneven distribution of the burdens and the benefits across regions and actors. Studies show that AI development reproduces structural inequalities between the Global South and the Global North (Mohamed et al., 2020; Trabelsi, 2024). The extreme market concentration reinforces this inequality: a small group of dominant platforms holds disproportionate control over behavioural and operational data, creating data monopolies and high barriers to entry for smaller actors (Zuboff, 2023; Schweitzer et al., 2019). By analysing the economic data value chain, Jia et al. (2025) argue that monetary value systematically accrues to aggregators and model developers, i.e., the dominant platforms, while data generators are largely excluded.
@@ 3. Methodology
To address the costs of hyper-datafication holistically, we estimate the energy use of large-scale data storage (environmental), analyse labour conditions using questionnaires administered to data workers and patterns of language representation in AI datasets (social), and draw on public data to examine economic costs associated with AI data (economical). This type of empirical analysis necessarily requires grounding in concrete cases and we therefore draw on a set of complementary data sources. First, the Hugging Face Hub serves as our primary lens on dataset growth trends. As one of the most widely used public repositories in AI development, it combines broad community adoption with detailed, consistent metadata, and has been used in meta-analyses in prior studies (Di Sipio et al., 2024; Castaño et al., 2024). Second, we use data workers in Kenya as a case study for data labour costs, as Kenya has been identified as a leading hub for AI-related gig work (World-Economic-Forum and Geneva-Graduate-Institute., 2025), frequently described as the “Silicon Savannah” (Graham and Mann, 2013). Finally, we draw on the IEA (International Energy Agency, 2025b) and Data Center Map (19) to contextualise the economic scale and distribution of data centre infrastructure. Together, these sources serve as strong indicators of broader trends, although each carries limitations that we discuss in Section 5.
Drawing on these sources, we aim to identify: (i) the environmental impacts of data storage under hyper-datafication; (ii) the actors and labour conditions that sustain hyper-datafication; (iii) the extent to which hyper-datafication represents society using language as a proxy; and (iv) the economic burdens associated with hyper-datafication.
@@ 3.1. Metadata of Hugging Face Datasets
Scope. We retrieved metadata on 1 December 2025, when the Hugging Face Hub hosted 570,802 datasets. Metadata extraction failed for 2.89% of repositories, yielding a final sample of 554,300 datasets. We collected repository-level metadata (identifiers, timestamps, downloads, and Hub-side storage), dataset-level metadata (dataset size), and contextual attributes (region, modality, task, and language). Hugging Face has options to query download statistics in two granularities: all time and last 30 days. As we retrieved our data on 1 December 2025, the statistics for last 30 days correspond to November 2025222Source code used to extract metadata from datasets hosted on the Hugging Face Hub is available at: https://github.com/saintslab/costs-of-hyperdatafication..
@@
Data collection. We collected individual repository-level metadata via the Hugging Face Hub Representational State Transfer (REST) Application Programming Interface (API) [^3] Using the endpoint https://huggingface.co/api/datasets/{id} (one request per dataset ID) and retrieving attributes via the API’s expand fields., a standard HyperText Transfer Protocol (HTTP) interface that returns structured metadata for a given dataset by its ID. Dataset size statistics were obtained separately from the Hugging Face Hub datasets-server API [^4] See https://github.com/huggingface/dataset-viewer for documentation of the datasets-server infrastructure., which computes dataset statistics and aggregates results across configurations and splits. Contextual attributes were extracted from the Hugging Face Hub Python API [^5] Accessing the public dataset catalogue at https://huggingface.co/datasets., a Python client library that wraps the Hub’s REST interface and enables programmatic iteration over the full public dataset catalogue. These attributes are primarily self-declared and therefore exhibit lower coverage. Appendix A.1 summarises all metadata attributes and coverage.
@@
Data analysis. We distinguish two storage measures: dataset size (the size of the raw data files in bytes) and Hub-side storage (the total space occupied on Hugging Face’s servers, including additional metadata and auxiliary files beyond the raw data). The former serves as a proxy for local storage footprint, and the latter for platform storage. Language tags were mapped to ISO-639 codes, an international standard for language classification. For the language analysis, dataset volumes were compared against global speaker distributions and web presence. As a proxy for the latter, we used page counts from Common Crawl, a publicly available web archive widely used for training language models. Extended descriptive statistics are provided in Appendix A.2.
3.2. Survey of Data Workers
@@
To assess social and labour costs, we conducted an online survey targeting data workers in Kenya, administered via a questionnaire.
@@
Questionnaire design and administration. The questionnaire was designed to address gaps in literature by examining critical aspects of data work, formulated as research questions covering demographics, working conditions, exposure to graphic content, and employment relationships with large technology companies. It was developed in accordance with established principles of questionnaire design (Roopa and Rani, 2012) and questions were kept concise and written in plain English to ensure accessibility. The questionnaire was developed by members of the research team and independently reviewed by others prior to deployment to ensure clarity and alignment with the study’s objectives. The final questionnaire comprised ten questions and is provided in Appendix B. It was administered via an online form and distributed through a messaging application to data workers who had voluntarily joined the channel organised by a local data workers’ collective. Data collection took place in December 2025.
Participants. The final sample includes 134 respondents located in Kenya. The sample comprises 57 females and 77 males, predominantly aged 20-40 years, with most reporting at least four years of experience. Respondents were informed about the study’s purpose and how their data would be used. They were not compensated, as institutional policy precludes participant payment. The survey was designed to take around five minutes to complete and the participants were free to participate and stop the survey at any time.
@@
Data analysis. We computed descriptive statistics and pairwise associations across working hours, experience, salary, exposure to graphic content, and employment type. Some questions allowed free-text responses, which, in a few cases, were used to harmonise categories. Free-text responses that could not be mapped consistently to the categories were excluded. Detailed preprocessing decisions and complete raw counts are reported in Appendix C.